OpenStudy (superman499):

During a construction project, engineers used explosives to excavate 140 feet of tunnel into a mountain. But because of time constraints and environmental concerns, they brought in a tunnel boring machine (TBM) to excavate the rest of the tunnel. The data table lists some observations an engineer made about the length of the tunnel after the TBM was introduced. How much of the tunnel was completed 26 days after the TBM was introduced? 1,170 feet 1,310 feet 1,260 feet 1,400 feet TBM Use (days) (feet) 6 410 15 815 20 1,040 28 1,400 33 1,625 38 1,850 45 2,165

OpenStudy (anonymous):

20 days =1040 +6/8 of 360 because from 20 to 28 days 360 was excavated 26 days= 1040 +270=1310